OpenStudy (anonymous):

1.Which property states that the order of numbers can be changed in an addition or multiplication problem? A. Associative B. Commutative C. Transitive D. Reflexive 2.Which property states that 1z = z? A. Reflexive B. Transitive C. Symmetric D. Multiplicative identity

OpenStudy (anonymous):

1. Commutative \(A+B = B+A\) and \( A\times B = B\times A \) 2. D. Multiplicative identity 1 is the multiplicative identity of \(\mathbb{Z} \).

OpenStudy (anonymous):

i thought it was associative for number 1

OpenStudy (anonymous):

can you explain to my why associative is not answer.

OpenStudy (anonymous):

me*